A toroidal sample of magnetic material of susceptibility x = 2 × 10^-2 is wound with 100 turns of wire carrying a current of 2 A. The toroid is 0.10 m long.

(a) Find the solenoidal current density.(2)

(b) Determine the magnetic field intensity H produced by the current.(2)

(c) Calculate μ, the magnetic permeability of the material.(2)

(d) Calculate the induced magnetization M in the material.(2)

(e) Calculate the magnetic field B resulting from the current and the magnetization of the

material


1
Expert's answer
2021-05-21T17:06:33-0400

Answer

A.

Current density

"j=\frac{I}{A}=\frac{2}{\pi (0.1) ^3}=63.69A/m^2"


B.

Magnetic field

"B=\frac{(1+x)NI}{2\pi R}"



"B=\frac{(1+0.2)(100)(2)}{2\pi (0.1)}=324.67T"


And

"H=\frac{B}{1+x}=\frac{324.67}{1+0.02}=318.3T"



C.

Magnetic permeability

"\mu=1+x=1+0.02=1.02"


D.

"M=Hx
\\M=318.4\times0.02=6.5"


E.

Magnetic field we have calculated previously

"B=324.67T"